# MERA Code: A Unified Framework for Evaluating Code Generation Across Tasks

### v2

@@ Appendix B Taxonomy details
Appendix C Evaluation results
@@
Model Private Score ruCodeEval CodeLinterEval ruCodeReviewer
@@
Gemini 2.5 flash 0.427 0.61 0.645 0.652 0.496 0.538 0.545 0.031 0.152 0.064 0.145 0.193
GPT-4.1 0.382 0.443 0.484 0.494 0.555 0.585 0.6 0.017 0.136 0.096 0.102 0.102
GPT-4o 0.381 0.529 0.559 0.567 0.479 0.518 0.527 0.018 0.133 0.078 0.094 0.1
GigaChat 2 Max 0.365 0.537 0.588 0.591 0.425 0.461 0.473 0.016 0.136 0.062 0.065 0.071
DeepSeek-Coder-V2-Inst 0.36 0.433 0.45 0.457 0.494 0.59 0.618 0.015 0.135 0.06 0.061 0.062
Seed-Coder-8B-Inst 0.345 0.317 0.317 0.317 0.655 0.655 0.655 0.019 0.156 0.035 0.052 0.06
@@
Qwen2.5-72B-Inst 0.254 0.174 0.177 0.177 0.481 0.497 0.5 0.023 0.158 0.048 0.104 0.136
Vikhr-YandexGPT-5-Lite-8B 0.187 0.035 0.041 0.043 0.407 0.515 0.518 0.015 0.126 0.022 0.023 0.023
Yi-Coder-9B-Chat 0.181 0.35 0.362 0.372 0.145 0.157 0.164 0.009 0.078 0.016 0.016 0.016
GigaCode 1.4 0.166 0.357 0.364 0.366 0.027 0.027 0.027 0.029 0.182 0.064 0.12 0.145
Mixtral-8x22B-Inst 0.028 0.0 0.0 0.0 0.027 0.045 0.045 0.017 0.135 0.016 0.025 0.025
@@
Model Total Score
@@
acc
GPT-4.1 0.377 0.144 0.034 0.297 0.418 0.162 0.344 0.639 0.45 0.48 0.494 0.416 0.66
GPT-4o 0.377 0.149 0.038 0.275 0.324 0.153 0.37 0.705 0.537 0.558 0.561 0.399 0.666
Gemini 2.5 flash 0.356 0.12 0.029 0.217 0.388 0.174 0.211 0.502 0.604 0.654 0.665 0.386 0.404
DeepSeek-Coder-V2-Inst 0.347 0.149 0.034 0.2 0.363 0.153 0.269 0.581 0.392 0.411 0.415 0.383 0.714
GigaChat 2 Max 0.346 0.106 0.014 0.294 0.332 0.175 0.137 0.396 0.53 0.57 0.579 0.342 0.68
Qwen2.5-Coder-32B-Inst 0.296 0.111 0.019 0.213 0.323 0.168 0.22 0.529 0.289 0.293 0.293 0.383 0.519
GigaCode 1.4 0.289 0.135 0.038 0.276 0.322 0.189 0.313 0.639 0.305 0.305 0.305 0.352 0.676
Qwen2.5-72B-Inst 0.285 0.144 0.024 0.252 0.362 0.128 0.189 0.476 0.157 0.163 0.165 0.349 0.702
Seed-Coder-8B-Inst 0.268 0.106 0.01 0.237 0.106 0.128 0.264 0.643 0.21 0.219 0.22 0.305 0.403
Yi-Coder-9B-Chat 0.203 0.135 0.024 0.192 0.067 0.138 0.229 0.59 0.173 0.197 0.201 0.252 0.364
Mixtral-8x22B-Inst 0.179 0.106 0.019 0.152 0.18 0.159 0.229 0.502 0.0 0.0 0.0 0.366 0.597
Vikhr-YandexGPT-5-Lite-8B 0.168 0.091 0.01 0.138 0.032 0.106 0.123 0.405 0.024 0.027 0.03 0.201 0.464

### v3

@@ Appendix D Taxonomy details
Appendix E Evaluation results
@@
Model
Private
Score
ruCodeEval CodeLinterEval ruCodeReviewer
@@
Gemini 2.5 flash 0.427 0.610 0.645 0.652 0.496 0.538 0.545 0.031 0.152 0.064 0.145 0.193
GPT-4.1 0.382 0.443 0.484 0.494 0.555 0.585 0.600 0.017 0.136 0.096 0.102 0.102
RuadaptQwen3-8B-Hybrid 0.380 0.419 0.548 0.598 0.500 0.537 0.545 0.034 0.181 0.042 0.078 0.118
Kimi-K2-Inst-0905 0.380 0.448 0.552 0.579 0.423 0.548 0.573 0.022 0.132 0.074 0.129 0.144
Qwen3-Coder-30B-A3B-Inst 0.373 0.391 0.418 0.433 0.621 0.632 0.636 0.018 0.139 0.058 0.074 0.083
A-vibe 0.367 0.263 0.405 0.439 0.614 0.664 0.673 0.035 0.182 0.023 0.071 0.102
GigaChat-2-Max 0.361 0.534 0.581 0.591 0.425 0.458 0.464 0.016 0.139 0.055 0.055 0.060
DeepSeek-Coder-V2-Inst 0.360 0.433 0.450 0.457 0.494 0.590 0.618 0.015 0.135 0.060 0.061 0.062
Seed-Coder-8B-Inst 0.345 0.317 0.317 0.317 0.655 0.655 0.655 0.019 0.156 0.035 0.052 0.060
Qwen3-Coder-480B-A35B-Inst 0.329 0.463 0.466 0.467 0.430 0.445 0.464 0.018 0.139 0.058 0.074 0.083
CodeLlama70b-Inst 0.310 0.040 0.120 0.171 0.507 0.908 0.973 0.007 0.051 0.006 0.023 0.025
@@
GigaChat-2-Lite 0.269 0.309 0.395 0.427 0.316 0.407 0.455 0.012 0.094 0.025 0.026 0.026
DeepSeek-V3.1 0.259 0.345 0.417 0.433 0.230 0.314 0.355 0.018 0.133 0.065 0.086 0.093
Qwen2.5-72B-Inst 0.254 0.174 0.177 0.177 0.481 0.497 0.500 0.023 0.158 0.048 0.104 0.136
Kodify Nano 0.226 0.191 0.313 0.354 0.235 0.432 0.473 0.005 0.043 0.003 0.003 0.003
Yi-Coder-9B-Chat 0.181 0.350 0.362 0.372 0.145 0.157 0.164 0.009 0.078 0.016 0.016 0.016
CodeLlama7b-Inst 0.176 0.000 0.000 0.000 0.315 0.602 0.664 0.000 0.001 0.000 0.000 0.000
GigaCode 1.4 0.166 0.357 0.364 0.366 0.027 0.027 0.027 0.029 0.182 0.064 0.120 0.145
gpt-oss-120b 0.106 0.004 0.016 0.024 0.208 0.328 0.355 0.003 0.025 0.000 0.000 0.000
CodeLlama34b-Inst 0.097 0.000 0.000 0.000 0.144 0.271 0.327 0.016 0.127 0.023 0.023 0.023
CodeLlama13b-Inst 0.071 0.000 0.000 0.000 0.058 0.225 0.355 0.000 0.000 0.000 0.000 0.000
Mixtral-8x22B-Inst 0.028 0.000 0.000 0.000 0.027 0.045 0.045 0.017 0.135 0.016 0.025 0.025
gpt-oss-20b 0.019 0.000 0.000 0.000 0.044 0.061 0.064 0.001 0.006 0.001 0.001 0.001
DeepSeek-Coder-V2-Lite-Inst 0.018 0.000 0.000 0.000 0.000 0.000 0.000 0.019 0.146 0.022 0.038 0.044
@@
Model
Total
Score
@@
EM
GPT-4.1 0.381 0.144 0.034 0.297 0.418 0.201 0.344 0.639 0.450 0.480 0.494 0.416 0.660
Gemini 2.5 flash 0.360 0.120 0.029 0.217 0.388 0.212 0.211 0.502 0.604 0.654 0.665 0.386 0.404
DeepSeek-Coder-V2-Inst 0.350 0.149 0.034 0.200 0.364 0.188 0.269 0.581 0.392 0.411 0.415 0.386 0.714
GigaChat-2-Max 0.338 0.091 0.014 0.291 0.318 0.223 0.145 0.361 0.524 0.559 0.567 0.262 0.686
RuadaptQwen3-8B-Hybrid 0.326 0.115 0.014 0.244 0.191 0.218 0.093 0.326 0.427 0.570 0.604 0.245 0.744
Qwen3-Coder-480B-A35B-Inst 0.326 0.082 0.019 0.202 0.411 0.231 0.220 0.573 0.421 0.424 0.427 0.342 0.540
Kimi-K2-Instruct-0905 0.325 0.077 0.029 0.232 0.188 0.153 0.233 0.507 0.448 0.554 0.579 0.342 0.571
A-vibe 0.317 0.115 0.024 0.307 0.157 0.232 0.119 0.348 0.257 0.411 0.470 0.279 0.724
Qwen3-Coder-30B-A3B-Inst 0.301 0.072 0.019 0.202 0.188 0.231 0.189 0.441 0.350 0.368 0.378 0.309 0.540
Qwen2.5-Coder-32B-Inst 0.300 0.111 0.019 0.213 0.323 0.206 0.220 0.529 0.289 0.293 0.293 0.386 0.519
GigaCode 1.4 0.293 0.135 0.038 0.276 0.322 0.230 0.313 0.639 0.305 0.305 0.305 0.352 0.676
Qwen2.5-72B-Inst 0.288 0.144 0.024 0.252 0.362 0.160 0.189 0.476 0.157 0.163 0.165 0.349 0.702
Seed-Coder-8B-Inst 0.271 0.106 0.010 0.237 0.106 0.161 0.264 0.643 0.210 0.219 0.220 0.305 0.403
DeepSeek-V3.1 0.248 0.106 0.024 0.293 0.084 0.159 0.189 0.520 0.282 0.372 0.390 0.275 0.372
GigaChat-2-Lite 0.229 0.048 0.005 0.144 0.099 0.107 0.035 0.278 0.251 0.344 0.372 0.148 0.712
Yi-Coder-9B-Chat 0.206 0.135 0.024 0.192 0.067 0.168 0.229 0.590 0.173 0.197 0.201 0.255 0.364
Mixtral-8x22B-Inst 0.182 0.106 0.019 0.152 0.180 0.194 0.229 0.502 0.000 0.000 0.000 0.366 0.597
Kodify Nano 0.181 0.082 0.000 0.092 0.041 0.112 0.053 0.300 0.173 0.299 0.341 0.104 0.475
CodeLlama70b-Inst 0.159 0.120 0.038 0.043 0.039 0.035 0.167 0.348 0.035 0.116 0.165 0.232 0.025
gpt-oss-120b 0.155 0.096 0.019 0.055 0.350 0.107 0.273 0.511 0.012 0.024 0.024 0.309 0.093
DeepSeek-Coder-V2-Lite-Inst 0.146 0.072 0.005 0.189 0.096 0.127 0.137 0.471 0.000 0.000 0.000 0.255 0.539
CodeLlama34b-Inst 0.141 0.120 0.029 0.120 0.019 0.176 0.088 0.317 0.000 0.000 0.000 0.195 0.472
CodeLlama7b-Inst 0.140 0.096 0.014 0.098 0.000 0.106 0.040 0.251 0.000 0.000 0.000 0.151 0.455
CodeLlama13b-Inst 0.127 0.082 0.024 0.143 0.016 0.156 0.093 0.361 0.001 0.003 0.006 0.171 0.419
gpt-oss-20b 0.095 0.038 0.000 0.062 0.305 0.148 0.013 0.084 0.000 0.000 0.000 0.285 0.115